OpenStudy (anonymous):

(g + 4) – 3g = 1 + g

OpenStudy (anonymous):

I think it g=3/2 I really don't know

jhonyy9 (jhonyy9):

g+4-3g=1+g -3g=-3 g=1

OpenStudy (anonymous):

Help! 2r-(5r)+13+2r